# Unveiling the Significance of Intrinsic Sensitivity and Multiple Scattering in Speckle Metrology

### v1

@@
Unveiling the Significance of Intrinsic Sensitivity and Multiple Scattering in Speckle Metrology
@@
We find that the width of $SCF(\Delta\lambda)$ changes significantly when the range of $\Delta\lambda$ changes. This dependency is due to the presence of averaging over $\alpha$ ($\lambda$ in this case) in EQ. 1. On the other hand, $S(\Delta\lambda)$ is essentially independent of the range of $\Delta\lambda$. Similar behaviour is also seen with different experimental data and even simulated speckles. Therefore, if $SCF(\Delta\alpha)$ is employed, its computation should encompass a substantial range of variation, approximately ten times larger than the apparent HWHM of the curve. This variability causes difficulties for the reliable comparison of different experimental platforms. Researchers can consider using $S(\Delta\alpha)$ as a more robust metric, and if $SCF(\Delta\alpha)$ is used then the range of values of $\alpha$ used in the calculation must be provided. Note that in FIG. 2, $SCF(\Delta\alpha)$ is normalised to $SCF(0)$, as the amplitude also varies depending on the subset used. Also note that just as $SCF(\Delta\alpha)$ is obtained by averaging over pixels, $S(\Delta\alpha)$ is also obtained by averaging over different pairs of images separated by the same $\Delta\alpha$ within the subset used.
@@
$S=\left|{\displaystyle\frac{\displaystyle\Big{\langle}\boldsymbol{E}^{\dagger}_{0}T^{\dagger}T^{\prime}\boldsymbol{E}_{0}\Big{\rangle}}{\displaystyle\sqrt{\Big{\langle}\boldsymbol{E}^{\dagger}_{0}T^{\dagger}T^{\prime}\boldsymbol{E}_{0}\Big{\rangle}}\sqrt{\Big{\langle}\boldsymbol{E}^{\dagger}_{0}T^{\dagger}T^{\prime}\boldsymbol{E}_{0}\Big{\rangle}}}}\right|^{2}.$ (7)
Here the effect of the change in the scattering medium takes the form of the operator $T^{\dagger}T^{\prime}$. Although it is not immediately evident, the dependency of the speckle’s sensitivity to the properties of the scattering medium is contained in this $T^{\dagger}T^{\prime}$ operator.
This can be shown by expressing the field on the observation plane as the superposition of the fields coming from all possible paths through the scattering medium [47, 67]. In a discrete description, a path can be defined as a succession of discrete elements of the scattering medium between which light propagates in straight lines. In this description, it can be shown (see Supplementary Information for a simplified derivation) that the Similarity becomes
@@
The data underpinning this work will be available through the University of St Andrews Research Data Portal.

### v2

# Determining intrinsic sensitivity and the role of multiple scattering in speckle metrology

@@
Determining intrinsic sensitivity and the role of multiple scattering in speckle metrology
@@
We find that the width of $SCF(\Delta\lambda)$ changes significantly when the range of $\Delta\lambda$ changes. This dependency is due to the presence of averaging over $\alpha$ ($\lambda$ in this case) in EQ. 1. On the other hand, $S(\Delta\lambda)$ is essentially independent of the range of $\Delta\lambda$. Similar behaviour is also seen with different experimental data and even simulated speckles. Therefore, if $SCF(\Delta\alpha)$ is employed, its computation should encompass a substantial range of variation, approximately ten times larger than the apparent HWHM of the curve. This variability causes difficulties for the reliable comparison of different experimental platforms. Researchers can consider using $S(\Delta\alpha)$ as a more robust metric, and if $SCF(\Delta\alpha)$ is used then the range of values of $\alpha$ used in the calculation must be provided. Note that in FIG. 2, $SCF(\Delta\alpha)$ is normalised to $SCF(0)$, as the amplitude also varies depending on the subset used.
@@
$S=\left|{\displaystyle\frac{\displaystyle\Big{\langle}\boldsymbol{E}^{\dagger}_{0}T^{\dagger}T^{\prime}\boldsymbol{E}_{0}\Big{\rangle}}{\displaystyle\sqrt{\Big{\langle}\boldsymbol{E}^{\dagger}_{0}T^{\dagger}T\boldsymbol{E}_{0}\Big{\rangle}}\sqrt{\Big{\langle}\boldsymbol{E}^{\dagger}_{0}T^{\prime\dagger}T^{\prime}\boldsymbol{E}_{0}\Big{\rangle}}}}\right|^{2}.$ (7)
This can be made more explicit by expressing the field on the observation plane as the superposition of the fields coming from all possible paths through the scattering medium [47, 67]. In a discrete description, a path can be defined as a succession of discrete elements of the scattering medium between which light propagates in straight lines. In this description, it can be shown (see Supplementary Information for a simplified derivation) that the Similarity becomes
@@
The data underpinning this work will be available through the University of St Andrews Research Data Portal at https://doi.org/10.17630/bcb2bff1-0d20-454a-89c5-8fd64b9e8ff8.